OpenStudy (anonymous):

Write an equation in slope-intercept form for the line through the two points. (-3,6); (2,1)

OpenStudy (ranga):

Find the slope using the formula m = (y2 - y1) / (x2 - x1) Then the equation of a line in slope-intercept form is: y = mx + b Put the value of m found earlier. This line passes through the given two points. Put any one of the points into y = mx + ba and solve for b.

OpenStudy (anonymous):

Why is there an a in y=mx+b

OpenStudy (ranga):

Typo. Put any one of the points into y = mx + b and solve for b.